OpenStudy (anonymous):

solve sqrt 9x+67=x+5

hero (hero):

Yeah, you're kinda late

OpenStudy (anonymous):

that is not right

jhonyy9 (jhonyy9):

yes this not is right because the resulted equation will be like factored hence (x+7)(x-6) and hence those solutions are x= -7 and x=6

jhonyy9 (jhonyy9):

Hero how you have wrote there will result x2 -x -42 =0

jhonyy9 (jhonyy9):

nikkiem now is right ???

hero (hero):

You wrote the answer incorrect the first time as well jhonyy9

jhonyy9 (jhonyy9):

check it before

hero (hero):

9x+67 = x^2+10x + 25 x^2+ x - 42 = 0 x^2+7x-6x-42 = 0 x(x+7)-6(x+7) = 0 (x+7)(x-6)=0 x =-7,6

hero (hero):

jhonyy 9, I don't understand how you don't understand how I got x^2+x-42 = 0 There's no way to solve this without getting there.

jhonyy9 (jhonyy9):

yeah but firstly you have wrote x= -6 and 7

hero (hero):

You did too

hero (hero):

But now you have erased it